Note that CPI data reports index values rather than rates of inflation. Yet, it is the rate of inflation that is interesting to business managers and economists, not the index values themselves. In class, we examined Bureau of Labor Statistics (BLS) indexed CPI data for Urban All Items. Recall that CPI data is normally reported monthly. The month-end index values for the seven months Jan ‘19 – Jul ‘19 look like this:     

Jan    253.1

Feb   254.1

Mar   255.0

Apr   255.2

May 255.3

Jun    256.2

Jul     256.3

Using these data, answer the following questions: Please show the calculation in detail.

What was the annualized rate of price increase observed during each of the following recent periods?:

  1. a) The month of Jul? [0.5%]
  2. b) The three months May-Jul [1.7%]
  3. c) The six months Feb-Jul[2.5%]
Answer #1

a) ((Inflation index of July/ Inflation Index of Jun)^12) - 1 = ((256.3/256.2)^12) - 1 = 0.469% or 0.5%

b) ((Inflation index of July/ Inflation Index of April)^4) - 1 = ((256.3/255.2)^4) - 1 = 1.735% or 1.7%

c) ((Inflation index of July/ Inflation Index of Jan)^2) - 1 = ((256.3/253.1)^2) - 1 = 2.544% or 2.5%